OpenStudy (anonymous):

Is 1563 years the right answer? A certain radioactive isotope has a half-life of approximately 900 years. How many years would be required for a given amount of this isotope to decay to 70% of that amount.

OpenStudy (anonymous):

we start with the formula \[A=A_0(\frac{1}{2})^{\frac{t}{900}}\]

OpenStudy (anonymous):

that formula is because you know the half life is 900

OpenStudy (anonymous):

then whatever you started with you will have 70% = .7 of it when you are done so set \[.7=(\frac{1}{2})^{\frac{t}{900}}\] and solve for t

OpenStudy (anonymous):

take the log of both sides to get \[ln(.7)=\frac{t}{900} ln(.5)\]

OpenStudy (anonymous):

solve for t via \[t=\frac{900ln(.7)}{ln(.5)}\]

OpenStudy (anonymous):

i get 463 rounded but i could have made a mistake

OpenStudy (anonymous):

oh. that is an avail answer. i was way off..haha

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

ok. thanks